OpenStudy (gio333):

Explain how to write the equation of a line if you know its y-intercept and the equation of a line to which it is parallel.

OpenStudy (anonymous):

post a picture of the graph

OpenStudy (gio333):

theres no graph

OpenStudy (larseighner):

For a line to be parallel to a given line it must have the same slope and a different y-intercept (if it has the same y-intercept it is the same line, not a parallel line. Put the given line in slope-intercept form (y = mx + b). Rip off the slope m. Use that for m and the y-intercept b to make the equation of your new line.

OpenStudy (larseighner):

That last b of course means the slope intercept of the new line, not copied from the given line.

OpenStudy (anonymous):

a line is a linear function y=ax+b with a the slope. Other lines have different a's and b's. But when they are parallel to the first line they have the same a. When they are perpendicular they have 1/a in front of the x. They say it intersects with y: when they say that they mean there's a b in the equation. When it doesn't intersect y except through the intersection of x-y (which is point 0-0) the linear function is y=ax.

OpenStudy (triciaal):

The equation of the line for any point on the line (x, y) is given by y = mx + b where m is the slope and c is the y- intercept. Parallel lines have the same slope. so, given m and c we can write the eqn. In addition, the slope is the change in y values/ corresponding change in x values when it is known that the lines goes through the 2 points (x, y) any point and (x1, y1) a given point.